Separate 75 into two numbers so that the sum of the two numbers equals 75. Three times the smaller number is five more than the larger number.
Answer by JBarnum(2146)

x+y=75
3x=y+5
3x-5=y
x+3x-5=75
4x=80
x=20
75-20=55
y=55